# Universal dissipators for driven open quantum systems and the correction to linear response

### v2

@@ 2.1 Stochastic Hamiltonian description — classical noise
Here $\mathcal{L}^{\prime}(t)=\mathcal{L}^{\prime}_{d}(t)+\mathcal{L}^{\prime}_{s}(t)$ and we introduced the cumulant symbol, which means $\bigr{\langle}\bigr{\langle}\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t-t^{\prime})\bigr{\rangle}\bigr{\rangle}=\bigr{\langle}\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t-t^{\prime})\bigr{\rangle}-\bigr{\langle}\mathcal{L}^{\prime}(t)\bigr{\rangle}\bigr{\langle}\mathcal{L}^{\prime}(t-t^{\prime})\bigr{\rangle}\bigr{\rangle}$, moreover $\bigr{\langle}\bigr{\langle}\mathcal{L}^{\prime}(t)\bigr{\rangle}\bigr{\rangle}=\bigr{\langle}\mathcal{L}^{\prime}(t)\bigr{\rangle}$, as follows from the standard definition of cumulant averages in statistics [49].
@@
$\mathbb{K^{\prime}}_{c}^{\rm{III}}(t)=i\int_{0}^{t}\!\int_{0}^{t_{1}}\bigl{\langle}\bigl{\langle}\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t_{1})\mathcal{L}^{\prime}(t_{2})\bigr{\rangle}\bigr{\rangle}\,dt_{1}dt_{2}\,.$ (13)
@@
$\mathbb{K^{\prime}}_{c}^{\rm{III}}(t)=i\int_{0}^{t}\!\int_{0}^{t_{1}}\bigl{\langle}\mathcal{L^{\prime}}_{s}(t)\bigl{[}\mathcal{L^{\prime}}_{d}(t_{1}),\mathcal{L^{\prime}}_{s}(t_{2})\bigr{]}\bigr{\rangle}\,dt_{1}dt_{2}\,.$ (15)
@@ 2.2 Bath-system description — quantum noise
Given the form above for the projector $\mathcal{P}$ onto the equilibrium subspace, we write $\mathcal{P}\rho^{\prime}=\mathrm{Tr}_{B}(\rho^{\prime})\otimes\rho_{B}$. Furthermore, utilizing the linearity of the trace and $\mathrm{Tr}\bigl{[}\mathcal{L}^{\prime}_{BS}(t)\rho^{\prime}_{S}(t)\bigr{]}\otimes\rho_{B}=0$ and $\mathrm{Tr}\bigl{[}\mathcal{L}^{\prime}_{S}(t)\mathcal{L}^{\prime}_{BS}(t^{\prime})\rho^{\prime}_{S}(t)\bigr{]}\otimes\rho_{B}=\mathrm{Tr}\bigl{[}\mathcal{L}^{\prime}_{BS}(t)\mathcal{L}^{\prime}_{S}(t^{\prime})\rho^{\prime}_{S}(t)\bigr{]}\otimes\rho_{B}=0$, we obtain:
@@ 3 Qubit with parametric dephasing
$\mathbb{K^{\prime}}^{\rm{III}}(t)=\frac{gD}{\tau}\int_{0}^{t}\int_{0}^{t-t^{\prime}}dt^{\prime}dt^{\prime\prime}\,c(t^{\prime}+t^{\prime\prime})\begin{bmatrix}0&&0&&\sin[\varphi]\\ 0&&0&&\cos[\varphi]\\ 0&&0&&0\\ \end{bmatrix}\,$. (36)
@@ 4 Qubit in a quantum environment
where $\Lambda_{0}$ is an overall coupling strength and $z_{l}=\xi_{l}-i\Gamma_{l}/2$ are the poles of the spectral density located only in the lower half of the complex plane, i.e, $\Gamma_{l}>0$. The poles have respective residues $r_{l}$, such that $ir_{l}>0$. To see this, one can evaluate $c(0)=\sum_{k}|g_{k}|^{2}$ that is a real positive number. Therefore, from Eq. (37), $i\Lambda_{0}^{2}\sum_{l=1}^{N}r_{l}$ must be also real and positive. Then, $ir_{l}$ is a real number and we restrict ourselves to positive cases, so that the effective couplings $\eta_{l}$, introduced later in Eq. (38), of the PMs are real. This treatment can be extended straightforwardly to complex PM coupling constants when needed [76]. The PMs consist of $N$ Bosonic modes associated with the poles. They are described by the annihilation(creation) operators $b_{l}^{(\dagger)}$, whose energy $\xi_{l}$ and decay rates $\Gamma_{l}$ are determined by the real and imaginary parts of the poles of the correlation function $c(t)$, respectively. The exact master equation governing the dynamics of the combined state $\rho(t)$ of the driven qubit along with the PMs is [76]
@@ 5 Conclusions
Going beyond this approximation, we have derived the first correction in the drive strength, in what we call the third-order master equation, which is still second-order in the bath-system coupling and first order in the drive strength, and therefore must be included in the usual Bloch-Redfield theory of relaxation. We showed how this third-order term notably improves the agreement with exact solutions of simple scenarios. The improvement is particularly good for the dynamics of the coherences of the reduced state. Quite interestingly, the third order in the quantum noise case predicts a striking effect. In fact, the last term in Eq. (41) suggests that it may be possible to dissipatively engineer a target steady state (even one with non-zero purity) by coherent pulses. This unleashes unprecedented capabilities in coherent control. An immediate application can be the design of a protocol for a coherent Overhauser effect, meaning the partial (limited-fidelity) transfer of the quantum state from a controlled spin to another, uncontrolled, one, only through the interaction with a shared environment. The design of this protocol should be quite straightforward thanks to the simple form of equation (41). Therefore, mimicking already existing protocols for the incoherent Overhauser effect [43, 83, 84], i.e., the one involving only the transfer of the populations of the spins, might be sufficient. With similar protocols, one could engineer coherent pulses for producing a steady state with some quantum correlations in a collection of spins by controlling individual spins that are only virtually coupled [83, 84].
@@ Appendix B Explicit calculations with the projection operator
where we omitted the integration bounds $\{0\,,\,t\}$ in all integrals to ease the notation. In this way, we obtain the following master equation in the Schrödinger picture,

### v3

@@ 2.1 Stochastic Hamiltonian description — classical noise
Here $\mathcal{L}^{\prime}(t)=\mathcal{L}^{\prime}_{d}(t)+\mathcal{L}^{\prime}_{s}(t)$ and we introduced the cumulant symbol, which means $\bigr\langle\bigr\langle\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t-t^{\prime})\bigr\rangle\bigr\rangle=\bigr\langle\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t-t^{\prime})\bigr\rangle-\bigr\langle\mathcal{L}^{\prime}(t)\bigr\rangle\bigr\langle\mathcal{L}^{\prime}(t-t^{\prime})\bigr\rangle$, moreover $\bigr\langle\bigr\langle\mathcal{L}^{\prime}(t)\bigr\rangle\bigr\rangle=\bigr\langle\mathcal{L}^{\prime}(t)\bigr\rangle$, as follows from the standard definition of cumulant averages in statistics [49].
@@
$\mathbb{K^{\prime}}_{c}^{\rm{III}}(t)=i\int_{0}^{t}\!\int_{0}^{t_{1}}\bigl\langle\bigl\langle\mathcal{L}^{\prime}(t)\mathcal{L}^{\prime}(t_{1})\mathcal{L}^{\prime}(t_{2})\bigr\rangle\bigr\rangle\,dt_{2}dt_{1}\,.$ (13)
@@
$\mathbb{K^{\prime}}_{c}^{\rm{III}}(t)=i\int_{0}^{t}\!\int_{0}^{t_{1}}\bigl\langle\mathcal{L^{\prime}}_{s}(t)\bigl[\mathcal{L^{\prime}}_{d}(t_{1}),\mathcal{L^{\prime}}_{s}(t_{2})\bigr]\bigr\rangle\,dt_{2}dt_{1}\,.$ (15)
@@ 2.2 Bath-system description — quantum noise
Given the form above for the projector $\mathcal{P}$ onto the equilibrium subspace, we write $\mathcal{P}\rho=\mathrm{Tr}_{B}(\rho)\otimes\rho_{B}$. Furthermore, utilizing the linearity of the trace and $\mathrm{Tr}\bigl[\mathcal{L}_{BS}(t)\rho_{S}(t)\bigr]\otimes\rho_{B}=0$ and $\mathrm{Tr}\bigl[\mathcal{L}_{D}(t)\mathcal{L}_{BS}(t-t^{\prime})\rho_{S}(t)\bigr]\otimes\rho_{B}=\mathrm{Tr}\bigl[\mathcal{L}_{BS}(t)\mathcal{L}_{D}(t-t^{\prime})\rho_{S}(t)\bigr]\otimes\rho_{B}=0$, we obtain:
@@ 3 Qubit with parametric dephasing
$\mathbb{K^{\prime}}^{\rm{III}}(t)=\frac{gD}{\tau}\int_{0}^{t}dt^{\prime}\int_{0}^{t-t^{\prime}}dt^{\prime\prime}\,c(t^{\prime}+t^{\prime\prime})\begin{bmatrix}0&&0&&\sin[\varphi]\\ 0&&0&&\cos[\varphi]\\ 0&&0&&0\\ \end{bmatrix}\,$. (36)
@@ 4 Qubit in a quantum environment
where $\Lambda_{0}$ is an overall coupling strength and $z_{l}=\xi_{l}-i\Gamma_{l}/2$ are the poles of the spectral density located only in the lower half of the complex plane, i.e, $\Gamma_{l}>0$. The poles have respective residues $r_{l}$, such that $ir_{l}>0$. To check this, one can evaluate $c(0)=\sum_{k}|g_{k}|^{2}$ that is a real positive number. Therefore, from Eq. (37), $i\Lambda_{0}^{2}\sum_{l=1}^{N}r_{l}$ must be also real and positive. Then, $ir_{l}$ is a real number and we restrict ourselves to positive cases, so that the effective couplings $\eta_{l}$, introduced later in Eq. (38), of the PMs are real. This treatment can be extended straightforwardly to complex PM coupling constants when needed [76]. The PMs consist of $N$ Bosonic modes associated with the poles. They are described by the annihilation(creation) operators $b_{l}^{(\dagger)}$, whose energy $\xi_{l}$ and decay rates $\Gamma_{l}$ are determined by the real and imaginary parts of the poles of the correlation function $c(t)$, respectively. The exact master equation governing the dynamics of the combined state $\rho(t)$ of the driven qubit along with the PMs is [76]
@@ 5 Conclusions
Going beyond this approximation, we have derived the first correction in the drive strength, in what we call the third-order master equation, which is still second-order in the bath-system coupling and first order in the drive strength, and therefore must be included in the usual Bloch-Redfield theory of relaxation. We showed how this third-order term notably improves the agreement with exact solutions of simple scenarios. The improvement is particularly good for the dynamics of the coherences of the reduced state. Quite interestingly, the third order in the quantum noise case predicts a striking effect. In fact, the last term in Eq. (41) suggests that it may be possible to dissipatively engineer a target steady state (even one with non-zero coherences) by coherent pulses. This unleashes unprecedented capabilities in coherent control. An immediate application can be the design of a protocol for a coherent Overhauser effect, meaning the partial (limited-fidelity) transfer of the quantum state from a controlled spin to another, uncontrolled, one, only through the interaction with a shared environment. The design of this protocol should be quite straightforward thanks to the simple form of equation (41). Therefore, mimicking already existing protocols for the incoherent Overhauser effect [43, 83, 84], i.e., the one involving only the transfer of the populations of the spins, might be sufficient. With similar protocols, one could engineer coherent pulses for producing a steady state with some quantum correlations in a collection of spins by controlling individual spins that are only virtually coupled [83, 84].
@@ Appendix B Explicit calculations with the projection operator
where the Liouvillean in the intermediate step is the full Liouvillean minus the bare part and we omitted the integration bounds $\{0\,,\,t\}$ in all integrals to ease the notation. In this way, we obtain the following master equation in the Schrödinger picture,